### v2

@@ 2.2 Theoretical Models
$\alpha_{1ik}$ $V_{ik}(1)-V_{ik}(0)$, the individual-specific ITT effect of $Z_{ik}$ on $V_{ik}$
$\beta_{D0ik}$ $\Pr(D_{ik}(0)=1)-\beta_{V0k}V_{ik}(0)$, the individual-specific probability of $D_{ik}(0)=1$ unexplained by $V_{ik}(0)$
$\beta_{D1ik}$ $\left\{\Pr(D_{ik}(1)=1)-\Pr(D_{ik}(0)=1)\right\}-\left\{\beta_{V1k}V_{ik}(1)-\beta_{V0k}V_{ik}(0)\right\}$, the individual-specific ITT effect of $Z_{ik}$ on $D_{ik}$ that is not transmitted through $V_{ik}$
$\gamma_{1ik}$ $\left\{Y_{ik}(1,0)-Y_{ik}(0,0)\right\}-\theta_{Vk}\left\{V_{ik}(1)-V_{ik}(0)\right\}$, the individual-specific direct effect of $Z_{ik}$ on $Y_{ik}$ that does not operate through $V_{ik}$ should the individual receive the control condition in the Phase II treatment
$\gamma_{2ik}$ $Y_{ik}(0,1)-Y_{ik}(0,0)$, the individual-specific direct effect of $D_{ik}$ on $Y_{ik}$ should the individual be assigned to the control condition in the Phase I treatment
$\gamma_{3ik}$ $\left\{Y_{ik}(1,1)-Y_{ik}(1,0)\right\}-\left\{Y_{ik}(0,1)-Y_{ik}(0,0)\right\}$, the individual-specific interaction effect of $Z_{ik}$ and $D_{ik}$ on $Y_{ik}$
$\alpha_{1k}$ $E\left(\alpha_{1ik}\mid S_{i}=k\right)$, the site-specific ITT effect of $Z_{ik}$ on $V_{ik}$
$\beta_{1k}$ $E\left\{D_{ik}(1)-D_{ik}(0)\mid S_{i}=k\right\}$, the site-specific ITT effect of $Z_{ik}$ on $D_{ik}$
$\beta_{2k}$ $E\left\{D_{ik}(1)\mid S_{i}=k\right\}$, the site-specific mean of $D_{ik}(1)$
$\gamma_{1k}$ $E\left(\gamma_{1ik}\mid S_{i}=k\right)$, the site-specific mean of $\gamma_{1ik}$
$\gamma_{2k}$ $E\left(\gamma_{2ik}\mid S_{i}=k\right)$, the site-specific mean of $\gamma_{2ik}$
$\gamma_{3k}$ $E\left(\gamma_{3ik}\mid S_{i}=k\right)$, the site-specific mean of $\gamma_{3ik}$
$\theta_{Vk}$ Site-specific linear association between $V_{ik}(z)$ and $Y_{ik}(z,d)$
Under the theoretical models (1), the cumulative ATE is the sum of four components as shown in equation (2):
$=E\left[E\left[\left\{\theta_{Y0ik}+\theta_{Vk}V_{ik}(1)+\gamma_{1ik}+\gamma_{2ik}+\gamma_{3ik}\right\}-\left\{\theta_{Y0ik}+\theta_{Vk}V_{ik}(0)\right\}\mid S_{i}=k\right]\right]$
$=\gamma_{1}+\gamma_{2}+\gamma_{3}+\theta_{V}\alpha_{1}.$ (2)
The last identity holds under the assumption that $\theta_{Vk}$ and $\alpha_{1k}$ are independent, which is one of the identification assumptions that we list below.
The cumulative ATE differs from the ITT effect in a two-phase study with noncompliance following Phase I. Specifically, the cumulative ATE represents the causal effect of the intended experimental treatment sequence compared to the control condition sequence, whereas the ITT effect reflects the causal effect of the initial treatment assignment regardless of the actual treatment sequences. The cumulative ATE also differs from the CATE or LATE, as it is defined for the entire targeted study population, while the CATE or LATE is specific to the subpopulation of compliers.
@@ 2.3 Identification Assumptions
Within-site zero covariance assumption. $\alpha_{0ik},\alpha_{1ik},\beta_{D0ik}$ and $\beta_{D1ik}$ are each uncorrelated with $\gamma_{\text{hik }}$ for $h=2,3$ within each site.
Assumption 2 implies the absence of individual-level pretreatment and posttreatment covariates that not only predict $V$ or $D$, or the individual-specific ITT effect of $Z$ on $V$ or $D$, but also predict the individual-specific causal effect of $D$ on $Y$ as a function of $Z$ within a site. Intuitively speaking, it assumes that the Phase I process is uncorrelated with the Phase II process within a site. The plausibility of this assumption increases when conditioning on observed pretreatment covariates $X$ and its interaction with $Z$ within each site. Assumption 2 differs from the conventional sequential ignorability assumption because, in the presence of omitted pretreatment covariatse $U$ that confounds the causal relationship between $D$ and $Y$, Assumption 2 still holds, whereas the sequential ignorability does not.
Violations of Assumption 2 are conceivable. For example, among students of the same level of kindergarten achievement, those who misbehave more due to attending a regular kindergarten class are perhaps more likely to change to a small class in Grade 1; moreover, such students may benefit more from attending a small class in Grade 1, in which case $\gamma_{2ik}$ and $\gamma_{3ik}$ would co-vary with $\beta_{D0ik}$. In another example, if students who achieve higher than others after attending a regular class in kindergarten are to benefit less from attending a small class in Grade 1 , then $\gamma_{2ik}$ would co-vary with $\alpha_{0ik}$.
@@
Between-site independence assumption. $\beta_{1k},\beta_{2k}$, and $\alpha_{1k}$ are each independent of $\theta_{Vk}$ and $\gamma_{hk}$ for $h=1,2,3$.
Assumption 3 implies the absence of site-level characteristics that predict not only the site-specific ITT effect of $Z$ on $V$, or the site-specific compliance rate, but also the site-specific association between $V$ and $Y$, or any other element of the site-specific cumulative ATE. Assumption 3 can be violated. For example, if schools with higher compliance rates in Grade 1 tend to derive a greater direct benefit of the kindergarten treatment on Grade 1 achievement, then $\beta_{1k}$ and $\gamma_{1k}$ are not independent.
@@ 2.4 Reduced-Form Models for the ITT Effects
Based on the theoretical models (1), we derive the reduced-form models for $V_{ik},D_{ik}$ and $Y_{ik}$ within each school, corresponding to which we can identify the school-specific ITT effects of $Z$ on $V,D$, and $Y$.
Reduced-form model for V. Because $E\left\{V_{ik}(z)\mid S_{i}=k\right\}=\alpha_{0k}+\alpha_{1k}z$, the school-specific ITT effect of $Z$ on $V$ is $E\left\{V_{ik}(1)-V_{ik}(0)\mid S_{i}=k\right\}=\alpha_{1k}$. The analytic model for obtaining the ITT effect of $Z$ on $V$ in school $k$ is given by
$V_{ik}=\left(1-Z_{ik}\right)V_{ik}(0)+Z_{ik}V_{ik}(1)=\alpha_{0k}+\alpha_{1k}Z_{ik}+e_{Vik},$ (3)
where $e_{Vik}=\alpha_{0ik}-\alpha_{0k}+\left(\alpha_{1ik}-\alpha_{1k}\right)Z_{ik}$ is an error term that, under Assumption 1, has mean zero and is uncorrelated with $Z_{ik}$ within school $k$, as we show in section S1.1 of the supplementary material.
Reduced-form model for $D$. Because $E\left\{D_{ik}(z)\mid S_{i}=k\right\}=\beta_{0k}+\beta_{1k}z$, where $\beta_{0k}=\beta_{D0k}+\alpha_{0k}\beta_{V0k}$ and $\beta_{1k}=\beta_{D1k}+\left(\alpha_{0k}+\alpha_{1k}\right)\beta_{V1k}-\alpha_{0k}\beta_{V0k}$, the school-specific ITT effect of $Z$ on $D$ is $E\left\{D_{ik}(1)-D_{ik}(0)\mid S_{i}=k\right\}=\beta_{1k}$. The analytic model for obtaining the school-specific ITT effect of $Z$ on $D$ in school $k$ is given by
$D_{ik}=\left(1-Z_{ik}\right)D_{ik}(0)+Z_{ik}D_{ik}(1)=\beta_{0k}+\beta_{1k}Z_{ik}+e_{Dik},$ (4)
where $e_{Dik}$ is an error term that, under Assumption 1, has mean zero and is uncorrelated with $Z_{ik}$ within school $k$, as we show in section S1.2 of the supplementary material.
Reduced-form model for $Y$. Under Assumption 2, we derive that
$E\left\{Y_{ik}(0)\mid S_{i}=k\right\}=\theta_{Y0k}+\theta_{Vk}\alpha_{0k}+\gamma_{2k}\beta_{0k}.$
$E\left\{Y_{ik}(1)\mid S_{i}=k\right\}=\theta_{Y0k}+\theta_{Vk}\left(\alpha_{0k}+\alpha_{1k}\right)+\gamma_{1k}+\left(\gamma_{2k}+\gamma_{3k}\right)\left(\beta_{0k}+\beta_{1k}\right).$
Hence, the school-specific ITT effect of $Z$ on $Y$ is
@@
where $\beta_{2k}=\beta_{0k}+\beta_{1k}$, which is the school-specific compliance rate under the Phase-I experimental condition. Note that $Y_{ik}(z)=Y_{ik}(z,D_{ik}(z))$. Section S1.3 of the supplementary material derives the analytic model for obtaining the ITT effect of $Z$ on $Y$ in school $k$ :
$Y_{ik}=\left(1-Z_{ik}\right)Y_{ik}\left(0,D_{ik}(0)\right)+Z_{ik}Y_{ik}\left(1,D_{ik}(1)\right)=\theta_{0k}+\theta_{1k}Z_{ik}+e_{Yik},$ (5)
where $\theta_{0k}=\theta_{\mathrm{Y}0k}+\theta_{Vk}\alpha_{0k}+\gamma_{2k}\beta_{0k};\theta_{1k}=\gamma_{1k}+\gamma_{2k}\beta_{1k}+\gamma_{3k}\beta_{2k}+\theta_{Vk}\alpha_{1k}$. Under Assumptions 1 and 2, the error term $e_{Yik}$ has mean zero and is uncorrelated with $Z_{ik}$ within school $k$.
School-level population model. We have derived that $\theta_{1k}$ is a function of $\alpha_{1k},\beta_{1k}$, and $\beta_{2k}$ in a population of schools. Let
@@
where $\epsilon_{k}=\left(\gamma_{1k}-\gamma_{1}\right)+\left(\gamma_{2k}-\gamma_{2}\right)\beta_{1k}+\left(\gamma_{3k}-\gamma_{3}\right)\beta_{2k}+\left(\theta_{Vk}-\theta_{V}\right)\alpha_{1k}$. We can show that, under Assumption 3, $\epsilon_{k}$ has mean zero and is uncorrelated with the predictors $\beta_{1k},\beta_{2k}$, and $\alpha_{1k}$ in model (6) when these predictors are each independent of the school-specific coefficients $\gamma_{1k},\gamma_{2k},\gamma_{3k}$, and $\theta_{Vk}$. After identifying $\alpha_{1k},\beta_{1k}$, and $\beta_{2k}$ for each school, we can subsequently obtain the values of $\gamma_{1},\gamma_{2},\gamma_{3}$, and $\theta_{V}$, and thereby identifying $\delta_{ATE}$ per Equation (2).
@@ 2.5 Estimation and Statistical Inference
Based on the above theoretical results, we propose a multisite IV analysis for estimating the cumulative ATE of a two-phase treatment sequence. The analysis involves two stages. In Stage 1, we estimate the ITT effects of $Z$ on $V,D$, and $Y$ within each school. In Stage 2, we obtain estimates of $\gamma_{1},\gamma_{2},\gamma_{3}$ and $\theta_{V}$ through a between-school analysis.
Stage 1 analysis. Through a school-by-school OLS analysis of the sample data with $Z_{ik}$ as the predictor of $V_{ik},D_{ik}$, and $Y_{ik}$ for individual $i$ in school $k$, we estimate the respective school-specific ITT effects ${\alpha}_{1k},{\beta}_{1k}$, and ${\theta}_{1k}$ as well as the intercept ${\beta}_{0k}$ and then compute $\hat{\beta}_{2k}=\hat{\beta}_{0k}+\hat{\beta}_{1k}$. These analytic models are consistent with the reduced-form models in equations (3), (4), and (5).
Stage 2 analysis. We then analyze via OLS a school-level regression model as specified in equation (6), replacing $\theta_{1k},\beta_{1k},\beta_{2k}$, and $\alpha_{1k}$ with their sample analogues obtained from the Stage 1 analysis. Corresponding to equation (2), the sample estimator of the cumulative ATE is
@@
MS2T-IV with covariance adjustment. Using the sample estimators $\hat{\alpha}_{1k},\hat{\beta}_{1k}$, and $\hat{\beta}_{2k}$ as predictors in the Stage 2 analysis may lead to finite-sample bias in the estimation of $\delta_{ATE}$. We recommend an additional adjustment for site mean-centered X and their potential interactions with $Z$ in the Stage 1 analysis. This additional adjustment is expected to reduce such finite-sample bias, increase estimation efficiency, and additionally increase the plausibility of Assumption 2 by conditioning on not just the observed time-varying covariates but also the observed pretreatment covariates. Let $\overline{\mathrm{X}}_{.k}$ denote the site mean in school $k$. Specifically, in the reduced-form models (3), (4), and (5), let $e_{Vik}=\alpha_{\mathrm{x}k}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)+\alpha_{\mathrm{x}zk}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)Z_{ik}+\varepsilon_{Vik},e_{Dik}=\beta_{\mathrm{x}k}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)+\beta_{\mathrm{x}zk}^{\prime}\left(\mathrm{X}_{ik}-\mathrm{X}_{.k}\right)Z_{ik}+\varepsilon_{Dik}$, and $e_{Yik}=\theta_{\mathrm{x}k}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)+\theta_{\mathrm{xz}k}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)Z_{ik}+\varepsilon_{Yik}$, respectively. The analytic models in the Stage 1 analysis then become:
@@
Because $\mathrm{X}_{ik}$ is site mean-centered, the new error terms $\varepsilon_{Vik},\varepsilon_{Dik}$, and $\varepsilon_{Yik}$ each have a mean of zero and are uncorrelated with $Z_{ik}$ under Assumption 1. The site-specific ITT effects $\alpha_{1k},\beta_{1k}$, and $\theta_{1k}$ remain unchanged; yet the efficiency of their sample estimators will improve due to the reduction in error variance. At Stage 2, we analyze the same school-level model as specified in equation (6) and obtain the sample estimator of the cumulative ATE in the same form as that in equation (7).
@@
$\begin{aligned} \operatorname{Var}\left(\hat{\delta}_{ATE}\right)=&\operatorname{Var}\left(\hat{\gamma}_{1}+\hat{\gamma}_{2}+\hat{\gamma}_{3}+\hat{\alpha}_{1}\hat{\theta}_{V}\right)\\ =&\operatorname{Var}\left(\hat{\gamma}_{1}\right)+\operatorname{Var}\left(\hat{\gamma}_{2}\right)+\operatorname{Var}\left(\hat{\gamma}_{3}\right)\\ &+\operatorname{Var}\left(\hat{\alpha}_{1}\right)\operatorname{Var}\left(\hat{\theta}_{V}\right)+2\alpha_{1}\operatorname{Var}\left(\hat{\theta}_{V}\right)+2\theta_{V}\operatorname{Var}\left(\hat{\alpha}_{1}\right)\\ &+2\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\gamma}_{2}\right)+2\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\gamma}_{3}\right)+2\operatorname{Cov}\left(\hat{\gamma}_{2},\hat{\gamma}_{3}\right)\\ &+2\alpha_{1}\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\theta}_{V}\right)+2\theta_{V}\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\alpha}_{1}\right)+2\alpha_{1}\operatorname{Cov}\left(\hat{\gamma}_{2},\hat{\theta}_{V}\right)\\ &+2\theta_{V}\operatorname{Cov}\left(\hat{\gamma}_{2},\hat{\alpha}_{1}\right)+2\alpha_{1}\operatorname{Cov}\left(\hat{\gamma}_{3},\hat{\theta}_{V}\right)+2\theta_{V}\operatorname{Cov}\left(\hat{\gamma}_{3},\hat{\alpha}_{1}\right),\end{aligned}.$

### v3

@@ 2.2 Theoretical Model
$\alpha_{1ik}$ $V_{ik}(1)-V_{ik}(0)$, individual-specific ITT effect of $Z_{ik}$ on $V_{ik}$
$\beta_{0ik}$ $D_{ik}(0)$ that is individual-specific
$\beta_{1ik}$ $D_{ik}(1)-D_{ik}(0)$, individual-specific ITT effect of $Z_{ik}$ on $D_{ik}$
$\beta_{2ik}$ $D_{ik}(1)$ that is individual-specific
$\gamma_{1ik}$ $\left\{Y_{ik}(1,0)-Y_{ik}(0,0)\right\}-\theta_{Vk}\left\{V_{ik}(1)-V_{ik}(0)\right\}$, individual-specific direct effect of $Z_{ik}$ on $Y_{ik}$ that does not operate through $V_{ik}$ should the individual receive the control condition in the Phase II treatment
$\gamma_{2ik}$ $Y_{ik}(0,1)-Y_{ik}(0,0)$, individual-specific direct effect of $D_{ik}$ on $Y_{ik}$ should the individual be assigned to the control condition in the Phase I treatment
$\gamma_{3ik}$ $\left\{Y_{ik}(1,1)-Y_{ik}(1,0)\right\}-\left\{Y_{ik}(0,1)-Y_{ik}(0,0)\right\}$, individual-specific interaction effect of $Z_{ik}$ and $D_{ik}$ on $Y_{ik}$
$\theta_{Vik}$ individual-specific linear association between $V_{ik}(z)$ and $Y_{ik}(z,d)$
$\alpha_{1k}$ $E\left(\alpha_{1ik}\mid S_{i}=k\right)$, site-specific ITT effect of $Z_{ik}$ on $V_{ik}$
$\beta_{1k}$ $E\left\{D_{ik}(1)-D_{ik}(0)\mid S_{i}=k\right\}$, site-specific ITT effect of $Z_{ik}$ on $D_{ik}$
$\beta_{2k}$ $E\left\{D_{ik}(1)\mid S_{i}=k\right\}$, site-specific mean of $D_{ik}(1)$
$\gamma_{1k}$ $E\left(\gamma_{1ik}\mid S_{i}=k\right)$, site-specific mean of $\gamma_{1ik}$
$\gamma_{2k}$ $E\left(\gamma_{2ik}\mid S_{i}=k\right)$, site-specific mean of $\gamma_{2ik}$
$\gamma_{3k}$ $E\left(\gamma_{3ik}\mid S_{i}=k\right)$, site-specific mean of $\gamma_{3ik}$
$\theta_{Vk}$ $E\left(\theta_{vik}\mid S_{i}=k\right)$, site-specific mean of $\theta_{Vik}$
@@ 2.3 Identification Assumptions
Within-site zero covariance assumption. $\mathrm{Cov}\left\{\gamma_{2ik},\beta_{1ik}\right\}=0$, $\mathrm{Cov}\left\{\gamma_{3ik},\beta_{2ik}\right\}=0$, and $\mathrm{Cov}\left\{\theta_{vik},\alpha_{1ik}\right\}=0$.
Assumption 2 rules out individual-level pretreatment covariates that predict not just the individual-specific ITT effect of $Z$ on $D$ but also the individual-specific causal effect of $D$ on $Y$ under assignment to the control condition in Phase I. It likewise rules out covariates that predict not only the individual-specific compliance in Phase II after being assigned to the experimental condition in Phase I but also the individual-specific interaction effect of $Z$ and $D$ on $Y$. Moreover, it rules out covariates that predict not only the individual-specific ITT effect of $Z$ on $V$ but also the individual-specific linear association between $V$ and $Y$. Intuitively speaking, this assumption requires that the Phase I process is uncorrelated with the Phase II process within a site.
Violations of Assumption 2 are conceivable. For example, students who behave well due to attending a small kindergarten class are perhaps less likely to change to a regular class in Grade 1; moreover, such students may benefit more from attending a small class in Grade 1, in which case $\gamma_{3ik}$ would co-vary with $\beta_{2ik}$. In another example, if students who are more likely to comply with the treatment assignment are to benefit more from attending a small class in Grade 1 , then $\gamma_{2ik}$ would co-vary with $\beta_{1ik}$.
@@
Between-site independence assumption. $\alpha_{1k}$, $\beta_{1k}$, and $\beta_{2k}$ are jointly independent of $\theta_{Vk}$, $\gamma_{2k}$, and $\gamma_{3k}$; furthermore, $\alpha_{1k}$, $\beta_{1k}$, and $\beta_{2k}$ are each independent of $\gamma_{1k}$.
Assumption 3 implies the absence of site-level characteristics that predict not only the site-specific ITT effect of $Z$ on $V$, the site-specific ITT effect of $Z$ on $D$, or the site-specific compliance rate among individuals assigned to the experimental condition in Phase I, but also the site-specific association between $V$ and $Y$, or any other element of the site-specific cumulative ATE including $\gamma_{1k}$, $\gamma_{2k}$, and $\gamma_{3k}$. The joint independence further implies that, when $\alpha_{1ik}$, $\beta_{1ik}$, or $\beta_{2ik}$ covary within a site, such covariance is independent of $\theta_{Vk}$, $\gamma_{2k}$, and $\gamma_{3k}$; the former is also independent of the within-site covariance between $\theta_{Vik}$, $\gamma_{2ik}$, or $\gamma_{3ik}$; similarly, the within-site covariance of the latter type is independent of $\alpha_{1k}$, $\beta_{1k}$, and $\beta_{2k}$. Assumption 3 can be violated. For example, if schools with higher compliance rates in Grade 1 tend to derive a greater direct benefit of the kindergarten treatment on Grade 1 achievement, then $\beta_{1k}$ and $\gamma_{1k}$ are not independent.
Theorem 1.
Under the theoretical model (1) and when Assumptions 1, 2, and 3 hold, the cumulative ATE is the sum of four components as shown in equation (6):
$=E\left[E\left[\left\{\theta_{Y0ik}+\theta_{Vik}V_{ik}(1)+\gamma_{1ik}+\gamma_{2ik}+\gamma_{3ik}\right\}-\left\{\theta_{Y0ik}+\theta_{Vik}V_{ik}(0)\right\}\mid S_{i}=k\right]\right]$
$=\gamma_{1}+\gamma_{2}+\gamma_{3}+\theta_{V}\alpha_{1}.$ (2)
@@ 2.4 Reduced-Form Models for the ITT Effects
This section shows that, once we identify the site-specific ITT effects of $Z$ on $V,D$, and $Y$, we can subsequently identify the elements in the cumulative ATE. Below we derive the reduced-form models for $V_{ik},D_{ik}$ and $Y_{ik}$ within each site.
Reduced-form model for V. Because $E\left\{V_{ik}(z)\mid S_{i}=k\right\}=\alpha_{0k}+\alpha_{1k}z$, the site-specific ITT effect of $Z$ on $V$ is $E\left\{V_{ik}(1)-V_{ik}(0)\mid S_{i}=k\right\}=\alpha_{1k}$. Given that $Z_{ik}$ is randomized, the analytic model for obtaining the ITT effect of $Z$ on $V$ in site $k$ is given by
$V_{ik}=\alpha_{0k}+\alpha_{1k}Z_{ik}+e_{Vik},$ (3)
where $e_{Vik}$ is an error term that, under Assumption 1, has mean zero and is uncorrelated with $Z_{ik}$ within site $k$.
Reduced-form model for $D$. Because $E\left\{D_{ik}(z)\mid S_{i}=k\right\}=\beta_{0k}+\beta_{1k}z$ where $\beta_{0k}=\beta_{2k}-\beta_{1k}$, the site-specific ITT effect of $Z$ on $D$ is $E\left\{D_{ik}(1)-D_{ik}(0)\mid S_{i}=k\right\}=\beta_{1k}$. Given that $Z_{ik}$ is randomized, the analytic model for obtaining the site-specific ITT effect of $Z$ on $D$ in site $k$ is given by
$D_{ik}=\beta_{0k}+\beta_{1k}Z_{ik}+e_{Dik},$ (4)
where $e_{Dik}$ is an error term that, under Assumption 1, has mean zero and is uncorrelated with $Z_{ik}$ within site $k$.
Reduced-form model for $Y$. As we show in section S1.1 of the supplementary material, under Assumptions 1 and 2, the site-specific ITT effect of $Z$ on $Y$ is
@@
where $\beta_{2k}=\beta_{0k}+\beta_{1k}$. Given that $Z_{ik}$ is randomized, the analytic model for obtaining the ITT effect of $Z$ on $Y$ in site $k$ is given by:
$Y_{ik}=\theta_{0k}+\theta_{1k}Z_{ik}+e_{Yik},$ (5)
where $\theta_{0k}=E\left\{Y_{ik}(0)\mid S_{i}=k\right\};\theta_{1k}=\gamma_{1k}+\gamma_{2k}\beta_{1k}+\gamma_{3k}\beta_{2k}+\theta_{Vk}\alpha_{1k}$. The error term $e_{Yik}$, under Assumption 1, has mean zero and is uncorrelated with $Z_{ik}$ within site $k$.
Site-level population model. We have derived that $\theta_{1k}$ is a function of $\alpha_{1k},\beta_{1k}$, and $\beta_{2k}$ in a population of sites. Let
@@
where $\epsilon_{k}=\left(\gamma_{1k}-\gamma_{1}\right)+\left(\gamma_{2k}-\gamma_{2}\right)\beta_{1k}+\left(\gamma_{3k}-\gamma_{3}\right)\beta_{2k}+\left(\theta_{Vk}-\theta_{V}\right)\alpha_{1k}$. We show in section S1.2 of the supplementary material that, under Assumption 3, $\epsilon_{k}$ has mean zero and is uncorrelated with the predictors $\beta_{1k},\beta_{2k}$, and $\alpha_{1k}$ in model (6). After identifying $\alpha_{1k},\beta_{1k}$, and $\beta_{2k}$ for each site, we can subsequently obtain the values of $\gamma_{1},\gamma_{2},\gamma_{3}$, and $\theta_{V}$, and thereby identifying $\delta_{ATE}$ per Equation (2).
2.5 The Distinction Between the Cumulative ATE and the ITT Effect on the Outcome
The cumulative ATE differs from the ITT effect in a two-phase study with noncompliance following Phase I. Specifically, the cumulative ATE represents the causal effect of the intended experimental treatment sequence compared to the control condition sequence, whereas the ITT effect reflects the causal effect of the initial treatment assignment regardless of the actual treatment sequences. The difference between the ITT effect and the cumulative ATE can be expressed as
$\mathrm{E}(\theta_{1k})-\delta_{\mathrm{ATE}}$ $=(\gamma_{1}+\gamma_{2}\beta_{1}+\gamma_{3}\beta_{2}+\theta_{V}\alpha_{1})-(\gamma_{1}+\gamma_{2}+\gamma_{3}+\theta_{V}\alpha_{1})$
$=\gamma_{2}(\beta_{1}-1)+\gamma_{3}(\beta_{2}-1).$
As expected, this difference vanishes under full compliance where $\beta_{1}=\beta_{2}=1$. The cumulative ATE also differs from the CATE or LATE, as it is defined for the entire targeted study population, while the CATE or LATE is specific to the subpopulation of compliers.
@@ 2.6 Estimation and Statistical Inference
Based on the above theoretical results, we propose a multisite IV analysis for estimating the cumulative ATE of a two-phase treatment sequence. The analysis involves two stages. In Stage 1, we estimate the ITT effects of $Z$ on $V,D$, and $Y$ within each site. In Stage 2, we obtain estimates of $\gamma_{1},\gamma_{2},\gamma_{3}$ and $\theta_{V}$ through a between-site analysis.
Stage 1 analysis. Through a site-by-site OLS analysis of the sample data with $Z_{ik}$ as the predictor of $V_{ik},D_{ik}$, and $Y_{ik}$ for individual $i$ in site $k$, we estimate the respective site-specific ITT effects ${\alpha}_{1k},{\beta}_{1k}$, and ${\theta}_{1k}$ as well as the intercept ${\beta}_{0k}$ and then compute $\hat{\beta}_{2k}=\hat{\beta}_{0k}+\hat{\beta}_{1k}$. These analytic models are consistent with the reduced-form models in equations (3), (4), and (5).
Stage 2 analysis. We then analyze via OLS a site-level regression model as specified in equation (6), replacing $\theta_{1k},\beta_{1k},\beta_{2k}$, and $\alpha_{1k}$ with their sample analogues obtained from the Stage 1 analysis. Corresponding to equation (2), the sample estimator of the cumulative ATE is
@@
MS2T-IV with covariance adjustment. Using the sample estimators $\hat{\alpha}_{1k},\hat{\beta}_{1k}$, and $\hat{\beta}_{2k}$ as predictors in the Stage 2 analysis may lead to finite-sample bias in the estimation of $\delta_{ATE}$. We recommend an additional adjustment for site mean-centered $X$ and their potential interactions with $Z$ in the Stage 1 analysis. This additional adjustment is expected to reduce such finite-sample bias, and increase estimation efficiency. Let $\overline{\mathrm{X}}_{.k}$ denote the site mean of $X$ in site $k$. Specifically, in the reduced-form models (3), (4), and (5), let $e_{Vik}=\alpha_{\mathrm{x}k}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)+\alpha_{\mathrm{x}zk}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)Z_{ik}+\varepsilon_{Vik},e_{Dik}=\beta_{\mathrm{x}k}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)+\beta_{\mathrm{x}zk}^{\prime}\left(\mathrm{X}_{ik}-\mathrm{X}_{.k}\right)Z_{ik}+\varepsilon_{Dik}$, and $e_{Yik}=\theta_{\mathrm{x}k}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)+\theta_{\mathrm{xz}k}^{\prime}\left(\mathrm{X}_{ik}-\overline{\mathrm{X}}_{.k}\right)Z_{ik}+\varepsilon_{Yik}$, respectively. The analytic models in the Stage 1 analysis then become:
@@
Because $\mathrm{X}_{ik}$ is site mean-centered, the new error terms $\varepsilon_{Vik},\varepsilon_{Dik}$, and $\varepsilon_{Yik}$ each have a mean of zero and are uncorrelated with $Z_{ik}$ under Assumption 1. The site-specific ITT effects $\alpha_{1k},\beta_{1k}$, and $\theta_{1k}$ remain unchanged; yet the efficiency of their sample estimators will improve due to the reduction in error variance. At Stage 2, we analyze the same site-level model as specified in equation (6) and obtain the sample estimator of the cumulative ATE in the same form as that in equation (7).
@@
$\operatorname{Var}\left(\hat{\delta}_{ATE}\right)=$ $\operatorname{Var}\left(\hat{\gamma}_{1}+\hat{\gamma}_{2}+\hat{\gamma}_{3}+\hat{\alpha}_{1}\hat{\theta}_{V}\right)$
$=$ $\operatorname{Var}\left(\hat{\gamma}_{1}\right)+\operatorname{Var}\left(\hat{\gamma}_{2}\right)+\operatorname{Var}\left(\hat{\gamma}_{3}\right)$
$+\operatorname{Var}\left(\hat{\alpha}_{1}\right)\operatorname{Var}\left(\hat{\theta}_{V}\right)+2\alpha_{1}\operatorname{Var}\left(\hat{\theta}_{V}\right)+2\theta_{V}\operatorname{Var}\left(\hat{\alpha}_{1}\right)$
$+2\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\gamma}_{2}\right)+2\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\gamma}_{3}\right)+2\operatorname{Cov}\left(\hat{\gamma}_{2},\hat{\gamma}_{3}\right)$
$+2\alpha_{1}\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\theta}_{V}\right)+2\theta_{V}\operatorname{Cov}\left(\hat{\gamma}_{1},\hat{\alpha}_{1}\right)+2\alpha_{1}\operatorname{Cov}\left(\hat{\gamma}_{2},\hat{\theta}_{V}\right)$
$+2\theta_{V}\operatorname{Cov}\left(\hat{\gamma}_{2},\hat{\alpha}_{1}\right)+2\alpha_{1}\operatorname{Cov}\left(\hat{\gamma}_{3},\hat{\theta}_{V}\right)+2\theta_{V}\operatorname{Cov}\left(\hat{\gamma}_{3},\hat{\alpha}_{1}\right).$